OpenStudy (anonymous):

A ball is thrown horizontally from the top of a building 21.8 m high. The ball strikes the ground at a point 101 m from the base of the building. The acceleration of gravity is 9.8 m/s 2 . Find the time the ball is in motion.

OpenStudy (abhisar):

Have you tried it first yourself?

OpenStudy (anonymous):

yes iv tryed but it keeps telling e it wrong

OpenStudy (abhisar):

Can you temme how are you doing it so that i could correct you....

OpenStudy (anonymous):

i did t= square root of (21.8/9.80

OpenStudy (abhisar):

Since,, the ball is thrown horizontally. The vertical component of initial velocity will be zero....ryt?

OpenStudy (abhisar):

Time of flight in this case will be equal to the time taken by a ball to reach the ground from 21.8 m during a free fall. Use the equation \(\sf s=ut+\frac{1}{2}\times g\times t^2\) take s=21.8 u=0 (initial velocity is 0 in case of free fall) g=9.8\(\approx\)10 Calculate t

OpenStudy (anonymous):

ok thank you

OpenStudy (abhisar):

Welcome.